### v1

# GOVERN: Gradient Orientation Vote Ensemble for Multi-Teacher Reinforced Distillation

Abstract
Pre-trained language models have become an integral component of question-answering systems, achieving remarkable performance. For practical deployment, it is critical to carry out knowledge distillation to preserve high performance under computational constraints. In this paper, we address a key question: given the importance of unsupervised distillation for student performance, how does one effectively ensemble knowledge from multiple teachers at this stage without the guidance of ground-truth labels? We propose a novel algorithm, GOVERN, to tackle this issue. GOVERN has demonstrated significant improvements in both offline and online experiments. The proposed algorithm has been successfully deployed in a real-world commercial question-answering system.
1 Introduction
Traditional search engine aims at deliver relevant web pages to satisfy users’ question, while sometimes the single paragraph that answer the question might buried deep in a web page, it asks for a web-based Open domain Question Answering (OpenQA) system to find that needle-in-a-haystack information (e.g. Qu et al., 2021; Zhang et al., 2023).
BERT-liked pre-trained language models have achieved state-of-the-art performance in OpenQA (e.g. Zhang et al., 2021). However, due to computational costs, the direct application of these models in real-time search engines like Google is currently unfeasible. For instance, the top-performing models on the Natural Question dataset, R2-D2 (Fajcik et al., 2021) and UnitedQA (Cheng et al., 2021) come with 1.29 billion and 2.09 billion model parameters, respectively. Further complicating matters is the fact that ensemble methods, which can enhance performance, entail even greater computational overheads.
The distillation of knowledge from multiple teachers has emerged as a powerful technique for improving the performance and generalization of deep neural networks while reducing the computational cost. This two-stage training paradigm, which training large model with limited labeled data as teacher and then using it to generate soft label on large amount unlabeled data for the purpose of student training, was first proposed by Hinton et al. (2015). Since the knowledge from single teacher may be biased and inaccurate, ensemble distillation from multiple teachers was considered by previous works to achieve more promising performance (e.g. You et al., 2017; Fukuda et al., 2017a).
Several dynamic distillation methods were proposed to solve the problem that different teacher is good at different sample and low-quality teachers may mislead the student. e.g. Yuan et al. (2021) proposed a novel RL-based approach to dynamically assigns weights among teachers, Cai et al. (2022) ensembles multi-teacher logits supervised by human-annotated labels in an iterative way. But these dynamic teacher selection methods need supervision signal as guidance, that means they can not apply to unsupervised distillation which is the most important stage in distillation (Su et al., 2021).
In this paper, we propose an algorithm called Gradient Orientation Vote Ensemble Reinforced distillatioN (GOVERN) to do sample-wise dynamic teacher selection without the need of label guidance. In the stage of unsupervised distillation, we take the gradient orientation provided by teachers as vote and follow the majority rule to do sample-wise dynamic teacher selection. While in supervised distillation stage, we use GOVERN to improve confidence aware ensemble method proposed by Zhang et al. (2022) with the help of human label.
Our main contributions are summarized as follows:
  • We propose GOVERN to do sample-wise dynamic teacher selection without the need of label guidance. We also give a mathematical proof that GOVERN can perform better than mean ensemble. To the best of our knowledge, GOVERN is the first method which can find sample-wise high-quality teachers without label guidance.
  • We propose a distillation framework for industrial application which combines GOVERN into both of the unsupervised and supervised distillation stage.
  • We conduct extensive experiments to demonstrate the effectiveness of GOVERN. Results show that GOVERN is benefit in both distillation stage and can boost a real-world question answering system’s performance.
2 Answer Selection Task
In a web-based Open domain Question Answering (OpenQA) system, the primary objective is to select the relevant paragraphs $A_{q}={a_{i}}_{i=1}^{N}\subset P_{q}$ which can solve the custom’s question $q\in Q$, where $P_{q}$ is a collection of paragraphs obtained in web pages retrieved by search engine. A classic framework of this system is made up of two-stage modules including retriever and ranker, where both modules can be distilled down to a task of classifying the relevance between a question and an answer. Our work focus on improving the performance of classification model with the limit of model size.
The classification model assesses the relevance of a paragraph, denoted as $p$, to a specific question, denoted as $q$, by calculating the relevance score, $f(q,p;\theta)$. This scoring function, $f$, which is parameterized by $\theta$, symbolizes the degree of relevance between the question $q$ and the paragraph $p$. In practical application, a score threshold is established for the purpose of classification.
During training, the classification model is optimized by minimizing the loss over training data:
$\mathop{min}\limits_{\theta}\mathop{\Sigma}\limits_{q\in Q}\mathop{\Sigma}\limits_{p\in P_{q}}l(y_{p}^{q},f(q,p;\theta))$ (1)
where $l$ is the loss function such as cross-entropy loss, margin loss or MSE loss, and $y_{p}^{q}$ is the relevance label of q-p pair.
3 Methodology
We use multiple teachers ensemble distillation as the method to improving the performance of online model with the constriction of computational cost. Within a frequently employed Knowledge Distillation (KD) framework, a large teacher model, denoted as T, is meticulously pretrained or finetuned well ahead of time. The knowledge contained within the teacher model is subsequently transferred to a smaller student model, denoted as S, by minimizing the disparity between the two. This process can be mathematically formulated:
$\mathop{min}\limits_{\theta}\mathop{\Sigma}\limits_{x}l(f^{S}(x;\theta),f^{T}(x;\Theta))$ (2)
where x embodies the input sample, while $f^{S}(\cdot)$ and $f^{T}(\cdot)$ denote the scoring function of the teacher and student models respectively. Additionally, $L(\cdot)$ serves as a loss function that calculates the variation between the behaviors of the two models.
Specifically, we first utilize unsupervised distillation on a vast amount of task-specific, unlabeled data, followed by supervised distillation on the labeled data. The procedures of the distillation can be viewed in Figure 1.
Figure 1: Procedures of Gradient Orientation Vote Ensemble Reinforced Distillation
3.1 Unsupervised Distillation
Unsupervised distillation, performed on a substantial amount of task-specific and unlabeled data, is vital for enhancing the performance of the student network. However, due to the absence of supervised signals, the prevalent unsupervised ensemble distillation method resorts to mean-ensemble to amalgamate the abilities of multiple models (You et al., 2017). Other studies have employed a weighted approach whereby individual teacher models are assigned varying weights to accentuate the contribution of higher performing models to knowledge transfer (e.g. Fukuda et al., 2017b; Kwon et al., 2020; Du et al., 2020; Liu et al., 2020). Methods to determine these weighting coefficients encompass weighting based on experience, calculating the weights based on logistic regression model, latent factor or multi-objective optimization in the gradient space.
While these weighting methods do account for the performance differences among various teacher models, they employ a uniform weighting coefficient for all samples during the distillation process. This approach neglects the varying emphasis on each teacher’s abilities and their respective confidence levels regarding samples with differing distributions.
Here, we propose a novel unsupervised voting method called Gradient Orientation Vote Ensemble Reinforced distillatioN (GOVERN), which does not rely on any human-annotated signals and dynamically assigns different teachers to different samples. In the following, we will introduce the implementation of this unsupervised distillation method and then mathematically prove its superiority over the mean-ensemble method.
3.1.1 GOVERN
In unsupervised distillation using mean-ensemble, for a sample, the distilled-model calculates $logit_{0}$, and N teacher-models calculate $logit_{i}$ respectively $(1<=i<=N)$. The distillation loss is:
$Dist(logit_{0},Mean(logit_{1},...,logit_{N}))$ (3)
where $Dist$ is a distance metric function that can be selected from MSE, cross-entropy, etc.
We take each teacher’s gradient descent orientation into consideration while doing ensemble. Specifically, when $logit_{i}>logit_{0}$, the gradient of $Dist(logit_{0},logit_{i})$ calculated is greater than 0, otherwise it is less than 0, so the gradient descent orientation is noted as:
$Grad_{i}$ $=SIGN(gradient(logit_{0},logit_{i}))$
=$=\left\{\begin{array}[]{rcl}1&&{logit_{i}>logit_{0}}\\ 0&&{logit_{i}=logit_{0}}\\ -1&&{logit_{i}<logit_{0}}\end{array}\right.$ (7)
The voted result is calculated as the sum of $Grad_{i}$:
$\chi(sample)=\left\{\begin{array}[]{rcl}1&&{\Sigma_{i=1}^{N}{Grad_{i}}>0}\\ 0&&{\Sigma_{i=1}^{N}{Grad_{i}}=0}\\ -1&&{\Sigma_{i=1}^{N}{Grad_{i}}<0}\end{array}\right.$ (11)
Each teacher is considerate as a voter in this way, then the loss for unsupervised distillation is represented as below:
$W_{i}$ =$=\left\{\begin{array}[]{rcl}1&&{\chi*Grad_{i}\geq 0}\\ 0&&{\chi*Grad_{i}<0}\\ \end{array}\right.$ (14)
$\mathcal{L}_{UD}$ $=MSE(logit_{0},\frac{\sum_{i=1}^{N}W_{i}logit_{i}}{\sum_{i=1}^{N}{W_{i}}})$ (15)
that means, we restrict our approach to guiding the student model’s training under the current sample solely by utilizing the majority of teacher models with consistent gradient orientations.
3.1.2 Proof
In this section, we mathematically prove that the sample-wise dynamic weighting ensemble algorithm GOVERN is better than mean-ensemble.
First, let’s consider the positive samples. For a binary classification model with an accuracy rate of p, the probability of correct classification after each sampling follows a Bernoulli distribution. Thus, the expected classification accuracy of a single teacher is p, and the variance is p(1-p).
To simplify computation, we assume the performance of the N teachers is consistent, i.e., $p=p_{1}=...=p_{N}$, where $p_{i}$ is the accuracy of $T_{i}$. The mean ensemble of N teachers is formulated as:
$X_{ME}=\frac{\Sigma_{i}X_{i}}{N}$ (16)
given that $X_{i}$ which follows Bernoulli distribution are independent and identically distribute, we obtain the conclusion that $E(X_{ME})=p$, $D(X_{ME})=p(1-p)/N$.
Therefore, compared to a single model, the mean-ensemble demonstrates better robustness.
Next, we consider the case where N teachers form a vote-ensemble classifier based on the principle of maximum voting. Then the expectation of the classifier is as follows:
$p_{0}=\sum_{m=\frac{N+1}{2}}^{N}C_{N}^{m}p^{m}(1-p)^{N-m}$ (17)
Utilizing mathematical induction, it is trivial to prove when $p>1/2$, $p_{0}>p$. This is called Condorcet’s jury theorem and details of proof can be found in (Sancho, 2022).
In the stage of distillation, with the assumption that teacher’s performance is better than student, we have $P(logit_{i}>logit_{0})>1/2$ for positive samples, where $logit_{i}$ is the output of $T_{i}$ and $logit_{0}$ is the output of student. So GOVERN is better than mean-ensemble.
For the negative samples, we can prove GOVERN is better in the same way.
3.2 Supervised Distillation: GOVERN-CA
Inspired by Confidence-Aware Multi-teacher Knowledge Distillation (CA-MKD) proposed by Zhang et al. (2022), we further develop GOVERN algorithm with the help of human label. On each training sample, we select the teachers which share the same gradient descent orientation with the human label. Furthermore, we assign weights among these selected teachers to reflect their sample-wise confidence by calculating the cross entropy loss between the prediction of teachers and human label:
$y(sample)=\left\{\begin{array}[]{rcl}1,&&{if\quad positive}\\ -1,&&{if\quad negative}\\ \end{array}\right.$ (20)
$W_{i}=\left\{\begin{array}[]{rcl}1&&{y*Grad_{i}>0}\\ 0&&{y*Grad_{i}\leq 0}\\ \end{array}\right.$ (23)
$\omega_{i}=\frac{W_{i}}{\Sigma_{j}W_{j}}(1-\frac{exp(L_{CE}^{i})}{\Sigma_{j}{W_{j}}exp(L_{CE}^{j})})$ (24)
where $L^{i}_{CE}$ denotes the cross entropy loss between the prediction of $i$-th teacher and human label, $Grad_{i}$ is defined in (11). The loss for supervised distillation is aggregated with calculated weights:
$\mathcal{L}_{SD}$ $=MSE(logit_{0},\Sigma_{i=1}\omega_{i}logit_{i})$ (25)
According to the aforementioned formulas, we only select teacher with the correct gradient descent orientation. Besides, the teacher whose prediction closely align with the ground-truth labels is assigned a greater weight $\omega_{i}$. This weighting is attributed to the model’s substantial confidence in making accurate judgments, thereby providing correct guidance.
Dataset #Question #Question-Paragraph Pair
unlabeled data 3,126,132 100M
train data 190,211 2,472,749
test data 3,301 93,446
Table 1: Dataset Statistic
Model Architecture Results
$n_{params}$ $n_{layers}$ $d_{model}$ $n_{heads}$ q R@P=90% qp R@P=90%
Teacher1-125M 125M 12 768 12 79.51% 70.52%
Teacher2-350M 350M 24 1024 16 81.79% 73.92%
Teacher3-1.5B 1.5B 48 1536 24 82.55% 73.09%
Teacher4-10B 10B 48 4096 64 83.06% 73.31%
Ensemble Model
Mean Ensemble - - - - 84.16% 76.71%
Logistic Regression Weighted Ensemble - - - - 83.44% 76.91%
Distilled Model
Mean Ensemble Distillation on unlabeled data 125M 12 768 12 82.04% 74.63%
LR Ensemble Distillation on unlabeled data 125M 12 768 12 81.98% 75.24%
GOVERN on unlabeled data 125M 12 768 12 83.65% 76.02%
+ CA-MKD on labeled data 125M 12 768 12 82.68% 75.67%
+ GOVERN-CA on labeled data 125M 12 768 12 83.69% 76.43%
Table 2: Results of offline experiments. Metrics denoted in bold represent the best results in the unsupervised distillation phase, while underscored and bolded denote the best results in the supervised distillation phase.
4 Experiments and Results
4.1 Dataset
The questions and relevant web-pages we use are collected from a commercial search engine, the objective is to select a paragraph which can answer the question from the web-pages. We set question-paragraph pairs as samples need to be classified. Hundred millions of unlabeled pairs are collected for unsupervised distillation, and we obtained millions of labeled pairs which are used for teacher’s fine-tune through crowd-sourcing annotators. The statistic of dataset is summarized in Table 1.
4.2 Experiment Details
Teacher Architecture In order to obtain multiple models with different structure and ability, we use pretrained models with different layer and fine-tune them on different samplings of the total labeled data. The specific structural parameters for each teacher model can be found in Table 2. Each model has been trained using a sample of 90% of the total data for training purposes.
Student Architecture Considering the computing resources and time consuming, we use the 12-layer transformer structure as the model for online deployment.
In the training procedure, we use the Adam optimizer (Kingma and Ba, 2015) with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. For all teacher models, we set the learning rate as 2e-5, the batch size as 64, and the warm-up step as 1000. The maximum length of input text is set as 384 and cross-entropy is used as loss function. In the distillation stage, we set the warm-up step as 1000, the learning rate as 2e-5 and the batch size as 64. The maximum length of input text is set as 384 and MSE is used as loss function. The best checkpoint is picked according to the performance on dev-set.
4.3 Evaluation Metrics
The metrics we used for experimental evaluation are introduced as below.
Precision-Recall is a useful measure of success of prediction when the classes are very imbalanced. $P=T_{p}/(T_{p}+F_{p}),R=T_{p}/(T_{p}+F_{n})$ ,where $T_{p}$, $F_{p}$ and $F_{n}$ represent for the number of true positives, false positives and false negatives, respectively.
Different threshold of a classifier leads to different Precision-Recall, follow the need of online system, we take recall value where precision equals to 90% as evaluation metrics.
q R@P=90% This metric only takes the paragraph with highest predicted score among all candidates under given question into consideration. A question is noted as $T_{p}$ if the score of selected answer is higher than threshold and the label is positive, while $F_{p}$ means the score of selected answer is higher than threshold but the label is negative. If the score of selected answer is lower than threshold but it does exist a positive answer for this question, we note it as $F_{n}$. This question granularity metric follows the behavior of web-based OpenQA system since system only displays the best answer was found, so it can best imitate model’s performance in online system.
qp R@P=90% This metric takes every qp-pair sample into consideration so it can reflect model’s general ability to find answers.
We also conduct a comparison called Good or Same or Bad (GSB) evaluation between two systems by inviting professional annotators to estimate which system produced a greater answer for each given question (Zhao et al., 2011). The gain of a new system can be formulated as:
$\Delta_{GSB}$ =#$=\frac{\#Good-\#Bad}{\#Good+\#Same+\#Bad}$ (26)
where #Good (or #Bad) denotes the number of questions that the new (or base) system provides better answer and #Same denotes the number of questions that answer are equal in quality.
4.4 Main Results
The main results of distillation methods comparison are shown in Table 2, we also display the results of teachers and ensemble methods. The methods used in the offline comparison experiments include:
Mean Ensemble We simply average the output of all teachers as the final predict score.
Logistic Regression Weighted Ensemble We trained a logistic regression model based on the development set to determine the weighting coefficients, and use them to obtain a weighted ensemble model.
CA-MKD This an algorithm proposed by Du et al. (2020) which adaptively assigns sample-wise reliability for each teacher prediction with the help of ground-truth labels, with those teacher predictions close to one-hot labels assigned large weights.
4.5 Online Experiment
To investigate the effectiveness of our proposed method in the real production environment, we deploy the proposed model in a commercial search engine, and conduct online experiments for comparison of MED and GOVERN.
Random Tail
$\Delta_{GSB}$ +4.5% +7.75%
$G:S:B$ 27: 364: 9 39: 353: 8
Table 3: Results of online experiments.
In contrast to random questions, tail questions are defined as those with a search frequency of less than 10 times per week. Given that heterogeneous search questions adhere to long-tail distributions, these tail questions constitute a significant portion of the questions processed by the search engine. It is evident that the proposed method consistently enhances the performance of the online QA system.
4.6 Ablation Study
We executed a series of experiments to illustrate the applicability of our method under various conditions. Due to computational resource constraints, we employed a 12-layer transformer as the teacher model and a 4-layer transformer as the student model for ablation study. The training data was divided into ten folds, with each teacher model trained on nine out of the ten folds, resulting in ten distinct teacher models. For the distillation process, we utilized fifty million unlabeled samples. The epoch for training was set to one.
Number of Teachers The impact of varying the number of teachers is illustrated in Figure 2. Experimental results indicate that the GOVERN algorithm consistently improves as the number of teachers increases. In contrast, mean-ensemble methods reach a performance plateau relatively quickly.
Figure 2: The effect of the number of teachers.
Effect of Single Teacher We further investigate the impact of varying the performance of a single teacher, with results presented in Table 4. The findings suggest that the GOVERN algorithm has the capacity to effectively select high-performing teachers, while simultaneously disregarding the noise generated by less effective ones.
qp prAUC
GOVERN with 5-teachers 88.19%
replace one teacher with 10B model 89.03%
replace one teacher with 4-layer model 88.11%
Table 4: Effect of Single Teacher.
5 Related Work
Following the seminal work of Hinton et al. (2015), several studies have sought to develop advanced ensemble algorithms for distillation. We categorize these works into two groups based on their dependency on ground-truth labels.
Unsupervised Ensemble Distillation There are a few works focused on the ensemble method on unsupervised data Li and Wang, 2019; Sui et al., 2020, these works simply use the average prediction of multiple teachers as the distillation signal. Recently, Wu et al. (2022) and Iliopoulos et al. (2022) made efforts on knowledge distillation with unlabeled examples, but these studies primarily concentrate on dynamically assigning weight to unlabeled examples. These approaches do not address the issue of teachers specializing in varying sample distributions.
Supervised Ensemble Distillation Li et al. (2021) first explored the idea of dynamic knowledge distillation with the help of ground-truth label. Yuan et al. (2021) proposed a novel RL-based approach, which dynamically assigns weights to teacher models at instance level. Cai et al. (2022) proposed algorithm ensembles multi-teacher logits supervised by human-annotated labels in an iterative way. Zhang et al. (2022) introduced confidence-aware mechanism on both predictions and intermediate features for multi-teacher knowledge distillation.
6 Conclusion
In this paper, we present a novel algorithm, GOVERN, which dynamically selects teachers based on their gradient descent orientation. This algorithm does not require ground-truth labels, making it suitable for use during unsupervised distillation stages. Additionally, it can be integrated with existing supervised ensemble methods. The effectiveness of our method is affirmed through extensive experimentation.
Limitations
The proposed GOVERN algorithm does not currently account for the varying performance levels of different teachers. This could be a shortcoming as it may be beneficial to assign a higher weight to more competent teachers, even if they share the same gradient descent orientation as other selected teachers.